A reaction to produce “substitute natural gas” is described by: 4 CO (g) + 8 H2 (g) → 3 CH4 (g) + CO2 (g) + 2 H2O (l) Use the following information, as necessary, to estimate DH for the above reaction.

2 C (graphite) + O2 (g) → 2 CO (g)ΔfH= –221.0 kJ/mol

2 CO (g) + O2 (g) → 2 CO2(g)ΔrH= –566.0 kJ/mol

2 H2 (g) + O2 (g) → 2 H2O (l)ΔfH= –571.6 kJ/mol

C (graphite) + 2 H2 (g) → CH4(g)ΔfH= –74.8 kJ/mol
Answer #1

Lets number the reaction as 0, 1, 2, 3, 4 from top to bottom

required reaction should be written in terms of other reaction

This is Hess Law

required reaction can be written as:

reaction 0 = -1.5 * (reaction 1) +0.5 * (reaction 2) +1 * (reaction 3) +3 * (reaction 4)

So, ΔHo rxn for required reaction will be:

ΔHo rxn = -1.5 * ΔHo rxn(reaction 1) +0.5 * ΔHo rxn(reaction 2) +1 * ΔHo rxn(reaction 3) +3 * ΔHo rxn(reaction 4)

= -1.5 * (-221.0) +0.5 * (-566.0) +1 * (-571.6) +3 * (-74.8)

= -747.5 KJ

Answer: -747.5 KJ